# On Fairness of Task Arithmetic: The Role of Task Vectors

### v1

@@ 4 Experimental Setup
For FFT, the pretrained model was fine-tuned on the combined training data from all subgroups of the target attribute (gender or race). Evaluation was then performed on the test data from each corresponding subgroup, enabling fine-grained assessment of both performance and fairness.
For LoRA, we followed the same training and evaluation procedure as FFT. The rank of LoRA’s adaptation modules was set to 8, following Ding et al. (2024b).
@@
To investigate whether incorporating task vectors from underperforming subgroups can improve fairness without sacrificing overall performance, we first identified the lowest-performing subgroups within each attribute based on the average of DPD and EOD under the FFT setting. We excluded the "others" group from this analysis as it does not reflect the characteristics of any specific subgroup. This selection was informed by both our experimental results and those reported in Ding et al. (2024b), which showed consistent patterns. For gender, the worst-performing subgroups were men and women; for race, they were Asian and Native American. We constructed a new model variant by injecting a worst-performing subgroup task vector worst-performing subgroup task vector into the base fine-tuned model:
$\theta_{\text{new}}=\theta_{\text{SFT}}+\lambda(\theta_{\text{worst-performingsubgroup}}-\theta_{0})$
where $\lambda$ controls the strength of the task vector injection. We varied $\lambda$ from from 0.0 to 1.0 at 0.2 intervals to analyze the effect of this targeted addition on subgroup fairness metrics and overall accuracy.
@@ 5 Results
Figure 3: Varying the task arithmetic coefficient $\lambda$ and comparing against FFT (purple dashed) and LoRA (orange dashed) for macro-averaged accuracy (left), demographic parity difference (DPD, center), and equalized odds difference (EOD, right). Evaluated on the gender subset of the data. Higher accuracy is better, while lower DPD and EOD indicate improved fairness. As $\lambda$ changes, task arithmetic maintains competitive accuracy and can reduce fairness gaps relative to the baselines.
Overview.
@@
In Figure 2(a), we compare the performance of FFT, LoRA, and task addition across gender subgroups; Figure 2(b) presents results for race subgroups. For task addition, we selected $\lambda=0.8$ for gender, $\lambda=0.5$ for race, as it achieved the highest average training accuracy across three random seeds within the tested range $\lambda\in[0.0,1.0]$. These visualizations provide a direct comparison of subgroup-wise model behavior.
From the subgroup-level bar plots in Figure 2, we observe that accuracy remains consistently high and comparable across all three adaptation methods, regardless of subgroup.
However, fairness metrics (DPD and EOD) show notable variation across methods. The impact of task addition on fairness is not consistent. Compared to FFT, task addition improved fairness in 5 out of 7 gender subgroups and in 3 out of 8 race subgroups. No single method yielded the best fairness performance across all demographic groups.
Our experimental results closely align with the findings reported in Ding et al. (2024b), particularly regarding the performance of FFT and LoRA on macro-averaged accuracy, DPD, and EOD. This consistency across demographic categories and with prior literature reinforces the robustness of our observations and supports the reliability of our evaluation framework.
@@
Figure 3 illustrates the overall performance of FFT, LoRA, and task arithmetic as the scaling coefficients for task addition vary from 0.0 to 1.0. We observe how varying the task-arithmetic coefficient $\lambda$ impacts macro-averaged accuracy (left), demographic parity difference (DPD, center), and equalized odds difference (EOD, right) on a gender subset of the data. As $\lambda$ increases from 0.0 to 0.2, we observe a peak in accuracy, but this configuration yields higher DPD and EOD, indicating reduced fairness. Beyond $\lambda=0.3$, accuracy remains competitive compared to FFT and LoRA, while both DPD and EOD progressively decline, suggesting that fairness improves without severely compromising performance. Notably, these task addition curves stay consistently lower than FFT and LoRA in terms of DPD and EOD at higher $\lambda$ values. Overall, this ablation could indicate that tuning $\lambda$ provides a practical mechanism for balancing accuracy and fairness objectives, offering guidelines for practitioners who wish to fine-tune fairness outcomes while maintaining strong predictive performance.
@@
Figure 4: Heatmaps of Accuracy (left), DPD (center), and EOD (right) for gender (top) and race (bottom) subgroups under the baseline FFT model ($\lambda=0.0$) and with increasing $\lambda$ values from 0.2 to 1.0 in 0.2 increments. The task vector for Men was added on the gender subset (top), and the task vector for Women was added on the gender subset (bottom). Darker cells indicate higher values on each metric’s scale; for DPD/EOD, lower values are better.
@@
Figure 5: Impact of injecting both the Men and Women subgroup task vectors into the FFT model on the gender data subset. The plot illustrates how scaling coefficient $\lambda$ reduces DPD and EOD, outperforming the baseline FFT (blue dashed) and LoRA (orange dashed), with negligible impact on macro-averaged accuracy.
5.2 Mixed trends when adding worst-performing subgroup task vectors
@@
We generally observe that increasing the scaling coefficient $\lambda$ tends to improve overall accuracy, consistent with the trends observed in Figure 3. However, the impact on fairness metrics (DPD and EOD) is more variable. In the gender-based plots, for example, the Asian subgroup consistently achieves the highest accuracy and lowest DPD/EOD—highlighting a recurring tradeoff where performance gains for one group may exacerbate disparities for others. When the Women task vector is added (Figure 4(b)), accuracy improves for the Trans Women subgroups. However, fairness metrics for subgroups such as Men tend to worsen as the scaling coefficient $\lambda$ increases. In Figure 4(a), injecting the Men task vector improves performance for some subgroups, yet Women consistently show lower accuracy and do not see consistent fairness improvements at higher $\lambda$. Some groups (e.g., Other, Trans Men, Trans Women) begin with relatively poor fairness under FFT and show partial improvements with task vector addition. Still, these improvements are not universal—for example, the Other subgroup often retains high EOD values regardless of $\lambda$. Likewise, Native American accuracy remains mostly unchanged across $\lambda$, while fairness metrics can deteriorate when injecting task vectors for other groups.
Overall, while increasing $\lambda$ can improve both accuracy and fairness for certain subgroups, these effects are not consistent across all configurations.
To visualize these results in more detail, Figure 5(a) shows macro-averaged accuracy, DPD, and EOD for the Men task vector added to the FFT model. The plots illustrate how varying the scaling coefficient $\lambda$ impacts overall performance and fairness, highlighting the effects of subgroup-specific task injection. We can observe in Figure 5(a) that injecting the Men task vector into the FFT model results in a slight accuracy gain and a clear monotonic decrease in both DPD and EOD as $\lambda$ increases—indicating a favorable and consistent improvement in fairness on the gender subset.
However, Figure 5(b) and the additional plots in Figures 11 and 12 in Appendix C.2 show more varied patterns as seen on Figures 4(a) and 4(b). When injecting the Native American task vector (Figure 12), accuracy remains stable while fairness seems to decrease (increased DPD and EOD). Asian (Figure 11) shows the same behavior as injecting the Men task vector (Figure 5(a)), positive increase of fairness metrics as $\lambda$ increases.
Taken together, these results confirm that task vector injection can shift subgroup-wise fairness and performance, but its effects are highly group-specific. While some subgroups (e.g., Men, Asian) exhibit smooth fairness gains, others (e.g., Women) show erratic behavior. This reinforces our earlier conclusion: simple task addition is not a guaranteed to have positive fairness influence, and more targeted strategies are likely needed - yet the scaling coefficient does show some relevant influence in fairness metrics.
@@ Conclusion.
In this study, we investigated the impact of a task arithmetic approach using task vectors on fairness, in comparison to conventional FFT and LoRA methods. We conducted detailed experiments to assess how the task addition affects prediction accuracy and fairness metrics, including the DPD and EOD across various subgroups. The results indicate that, with appropriate settings of the scalar coefficient $\lambda$, the task arithmetic method can improve DPD and EOD without significantly compromising overall model accuracy. Notably, using low to moderate values of the task vector coefficient effectively reduced prediction bias in minority groups compared to FFT and LoRA.
Furthermore, the task arithmetic framework allows for subgroup-specific evaluation and adjustment of model updates, enhancing interpretability—a key advantage of this method in the context of fairness. This interpretability facilitates the mitigation of excessive bias or adverse effects on particular groups, ultimately enabling more balanced model training.
@@
Despite these promising results, several challenges remain. The effectiveness of task arithmetic depends on dataset characteristics and subgroup distributions, necessitating further investigation into its generalizability across different tasks and domains. Moreover, future work should explore algorithms for automatically optimizing the scalar coefficient $\lambda$ and for balancing trade-offs among multiple subgroups.
In summary, our study demonstrates that task arithmetic using task vectors offers a promising approach for controlling model fairness. Further experimental validation, application to diverse tasks, and developing trade-off optimization methods are essential for improving fairness in broader and more realistic deployment scenarios.
Appendix A Fairness metrics
A.1 Demographic Parity Difference (DPD) (Agarwal et al., 2018, 2019)

### v2

@@ 4.2 Protocol
For FFT, the pretrained model was fine-tuned on the combined training data from all subgroups of the target attribute (gender or race). Evaluation was then performed on the test data from each corresponding subgroup, enabling fine-grained assessment of both performance and fairness. For LoRA, we followed the same training and evaluation procedure as FFT. The rank of LoRA’s adaptation modules was set to 8, following Ding et al. (2024).
@@
To investigate whether incorporating task vectors from underperforming subgroups can improve fairness without sacrificing overall performance, we first identified the lowest-performing subgroups within each attribute based on the average of DPD and EOD under the FFT setting. We excluded the “others” group from this analysis as it does not reflect the characteristics of any specific subgroup. This selection was informed by both our experimental results and those reported in Ding et al. (2024), which showed consistent patterns. For gender, the worst-performing subgroups were men and women; for race, they were Asian and Native American. We constructed a new model variant by injecting a worst-performing subgroup task vector into the base fine-tuned model:
$\theta_{\text{new}}=\theta_{\text{SFT}}+\lambda(\theta_{\text{worst-performing subgroup}}-\theta_{0}),$
where $\lambda$ controls the strength of the task vector injection. We varied $\lambda$ from 0.0 to 1.0 at 0.2 intervals to analyze the effect of this targeted addition on subgroup fairness metrics and overall accuracy.
@@ 5 Results
5.1 Theoretical intuition.
To situate the subsequent analysis, we first provide a high-level rationale for why task vector scaling interacts systematically with group-wise fairness metrics. Scaling the task vectors modifies the model parameters in directions associated with each subgroup, and the resulting model $\theta(\bm{\lambda})$ can be viewed as a perturbation of the balanced reference point $\bar{\theta}$ obtained at $\lambda_{g}=1$ for each $g$. Since group disparities are minimized at this balanced model, the behavior of fairness metrics is largely governed by the magnitude of the deviation $\theta(\bm{\lambda})-\bar{\theta}$ and by the norms of the subgroup-specific task vectors that induce this deviation. For convenience, we write $\bm{\lambda}=(\lambda_{g})_{g=1}^{G}$ for the vector of subgroup-specific scaling coefficients, with the balanced setting corresponding to $\lambda_{g}=1$ for all $g$.
We complement our empirical findings with an analytical upper bound that links task vector scaling to fairness metrics. We make several assumptions for the analysis: [A1] the prediction score $p_{\theta}(x)$ is $L$-Lipschitz in $\theta$ and group-wise calibrated to its binarized prediction (see Appendix C), [A2] a task vector for each group $g$ (i.e. $\Delta\theta_{g}$) is obtained under an identical optimization protocol, [A3] the scaling coefficients satisfy the normalization $\sum_{g}\lambda_{g}=G$, [A4] the group distributions differ only through the sensitive attribute so that the balanced model $\bar{\theta}=\theta_{0}+\frac{1}{G}\sum_{g}\Delta\theta_{g}$ satisfies $\operatorname{DPD}(\bar{\theta})=0$, [A5] near the binarization threshold, both $p_{\theta}(X)$ and $p_{\bar{\theta}}(X)$ have uniformly bounded class-conditional densities, i.e., for some $B_{y}>0$ for each true label $y$ and sufficiently small $\gamma$.
Assumptions [A1] and [A5] impose mild regularity and calibration conditions on the prediction scores, whereas [A2][A4] correspond to comparatively natural requirements on the training setup. Under those assumptions, the DPD and EOD satisfy the following inequalities:
Theorem (informal).
Consider the merged model $\theta(\bm{\lambda})=\theta_{0}+\sum_{g}\lambda_{g}\,\Delta\theta_{g}$. Then DPD and EOD satisfy:
$\operatorname{DPD}(\theta(\bm{\lambda}))$ $\leq U(\bm{\lambda})=2L\sum_{g}\big|\lambda_{g}-1\big|\,\|\Delta\theta_{g}\|_{2},\quad\text{for a Lipschitz constant $L$,}$
$\mathrm{EOD}(\theta(\bm{\lambda}))$ $\leq\operatorname{EOD}(\bar{\theta})+4\sqrt{(B_{0}+B_{1})\,U(\bm{\lambda})}.$
A full derivation and tighter constants are provided in Appendix C (Proposition 1,2).
These upper bounds explain the empirical trends observed in Figure 3. Both DPD and EOD share a common upper-bound term $U(\bm{\lambda})$, which monotonically decreases as $\lambda~(=\lambda_{1}=\cdots=\lambda_{G})$ approaches $1$ for each $g$. Notably, the bound $\operatorname{DPD}(\theta(\bm{\lambda}))\leq U(\bm{\lambda})$ together with $U(\bm{\lambda})\to 0$ as $\lambda_{g}\to 1$ implies that deviations of $\bm{\lambda}$ exert a pronounced influence on DPD. $B_{0}$ and $B_{1}$ are class-specific constants that upper-bound the score density near the decision threshold for y=0 and y=1, respectively. Intuitively, this bound shows that deviations of the scaling coefficient $\bm{\lambda}$ from the balanced setting ($\lambda_{1}=\cdots=\lambda_{G}=1$) enlarge disparities in proportion to the norms of subgroup task vectors.
@@
Figure 3: Varying the task arithmetic coefficient $\lambda$ and comparing against FFT (dashed baseline blue line) and LoRA (orange dashed) for macro-averaged accuracy (left), demographic parity difference (DPD, center), and equalized odds difference (EOD, right) on the gender subset. Higher accuracy is better; lower DPD/EOD indicate improved group fairness. For $\lambda\gtrsim 0.3$, task addition maintains competitive accuracy while typically lowering DPD/EOD relative to both baselines.
5.2 Empirical results overview.
Key takeaway. Across all datasets and architectures we study, task arithmetic exposes two complementary mechanisms for navigating fairness–accuracy trade-offs. First, sweeping a single scalar coefficient $\lambda$ over uniformly merged subgroup task vectors yields a smooth fairness–utility frontier: over a broad range of $\lambda$, task addition matches FFT/LoRA (or SFT/LoRA) accuracy while substantially reducing DPD and EOD on Berkeley D‑Lab hate speech and Civil Comments, and in our additional vision experiments on UTKFace, across LLaMA‑2, DistilBERT, Qwen‑2.5, and ViT‑Base. Second, injecting individual subgroup task vectors into an FFT model (Section 5.4) produces structured, subgroup‑dependent shifts in fairness: some subgroup vectors (e.g., Men, Asian) move the model along favorable fairness–accuracy frontiers, while others (e.g., Native American) can worsen DPD/EOD even when accuracy remains stable. This behavior is consistent with the group‑weighted ERM view developed in Section 5.1 and Appendix C.2, where sensitivity to $\lambda$ scales with $\|\Delta\theta_{g}\|_{2}$. Together, these observations show that task arithmetic provides both a global control knob (a uniform $\lambda$) and subgroup‑targeted knobs (the $\Delta\theta_{g}$) that are not exposed by standard FFT/LoRA training. The empirical trends and their first‑order mechanism also align with prior literature: our macro‑averaged accuracy, DPD, and EOD findings for FFT and LoRA are consistent with (Ding et al., 2024).
@@
Figures 2(a) and 2(b) instantiate this story for hate‑speech detection on Berkeley D‑Lab with LLaMA‑2 by comparing FFT, LoRA, and task addition across gender and race subgroups. For task addition, we selected $\lambda=0.8$ for gender and $\lambda=0.5$ for race, as these achieved the highest average training accuracy across three random seeds within the tested range $\lambda\in[0.0,1.0]$. These visualizations provide a direct comparison of subgroup‑wise model behavior. From the subgroup‑level bar plots in Figure 2, we observe that accuracy remains consistently high and comparable across all three adaptation methods, regardless of subgroup. On Civil Comments, on both DistilBERT and Qwen‑2.5, task addition similarly reduces group disparities while keeping accuracy competitive (see Appendix F and Table 5 for full confidence intervals and results).
We also observe that, relative to FFT, task addition improves fairness in five of seven gender subgroups and in three of eight race subgroups, with no single method dominating across all groups. The subgroup‑dependent nature of these effects motivates treating $\lambda$ as a deliberate tuning knob and inspecting subgroup behavior explicitly. As shown in Appendix C.2, task addition realizes a group‑weighted ERM in the linearized model: concretely, $\theta(\bm{\lambda})=\theta_{0}+\sum_{g}\lambda_{g}\,\Delta\theta_{g}$ coincides with the one‑step minimizer of a first‑order surrogate where subgroup $g$ is re‑weighted by $\lambda_{g}$. This explains the smooth fairness–utility frontier traced by sweeping $\lambda$, and Theorem Theorem (informal) predicts larger parity swings for groups with larger $\|\Delta\theta_{g}\|_{2}$. The observed curves in Fig. 3 align with those predictions without further assumptions.
@@
Figure 3 illustrates the overall performance of FFT, LoRA, and task arithmetic as the scaling coefficients for task addition vary from 0.0 to 1.0. We observe how varying the task‐arithmetic coefficient $\lambda$ impacts macro-averaged accuracy (left), demographic parity difference (DPD, center), and equalized odds difference (EOD, right) on a gender subset of the data. As $\lambda$ increases from 0.0 to 0.2, we observe a peak in accuracy, but this configuration yields higher DPD and EOD, indicating reduced fairness. Beyond $\lambda=0.3$, accuracy remains competitive compared to FFT and LoRA, while both DPD and EOD progressively decline, suggesting that fairness improves without severely compromising performance. Notably, these task addition curves stay consistently lower than FFT and LoRA in terms of DPD and EOD at higher $\lambda$ values. Overall, this ablation could indicate that tuning $\lambda$ provides a practical mechanism for balancing accuracy and fairness objectives, offering guidelines for practitioners who wish to fine‐tune fairness outcomes while maintaining strong predictive performance.
@@
Figure 4: Heatmaps of Accuracy (left), DPD (center), and EOD (right) for Men (top) and Women (bottom) subgroups under the baseline FFT model ($\lambda=0.0$) and with increasing $\lambda$ values from 0.2 to 1.0 in 0.2 increments. The task vector for Men was added on the gender subset (top), and the task vector for Women was added on the gender subset (bottom). Darker cells indicate higher values on each metric’s scale; for DPD/EOD, lower values are better.
@@
Figure 5: Impact of injecting the Men (a) and Women (b) subgroup task vectors into the FFT model on the gender data subset. The plot illustrates how scaling coefficient $\lambda$ reduces DPD and EOD, outperforming the baseline FFT (blue dashed) and LoRA (orange dashed), with negligible impact on macro-averaged accuracy.
5.4 Subgroup-targeted vectors: gains with trade-offs
@@
We generally observe that increasing the scaling coefficient $\lambda$ tends to improve overall accuracy, consistent with the trends observed in Figure 3. However, effects are not uniform across all subgroups. In the race-based plots, for example, the Asian subgroup consistently achieves the highest accuracy and lowest DPD/EOD—highlighting a recurring tradeoff where performance gains for one group may exacerbate disparities for others. When the Women task vector is added (Figure 4(b)), accuracy improves for the Trans Women subgroups. However, fairness metrics for subgroups such as Men tend to worsen as the scaling coefficient $\lambda$ increases.
In Figure 4(a), injecting the Men task vector improves performance for some subgroups, yet Women consistently show lower accuracy and do not see consistent fairness improvements at higher $\lambda$. Some groups (e.g., Other, Trans Men, Trans Women) begin with relatively poor fairness under FFT and show partial improvements with task vector addition. Still, these improvements are not universal—for example, the Other subgroup often retains high EOD values regardless of $\lambda$. Likewise, Native American accuracy remains mostly unchanged across $\lambda$, while fairness metrics can deteriorate when injecting task vectors for other groups. To visualize these results in more detail, Figure 5(a) shows macro-averaged accuracy, DPD, and EOD for the Men task vector added to the FFT model. The plots illustrate how varying the scaling coefficient $\lambda$ impacts overall performance and fairness, highlighting the effects of subgroup-specific task injection. We can observe in Figure 5(a) that injecting the Men task vector into the FFT model results in a slight accuracy gain and a clear monotonic decrease in both DPD and EOD as $\lambda$ increases—indicating a favorable and consistent improvement in fairness on the gender subset.
However, Figure 5(b) and the additional plots in Figures 11 and 12 in Appendix E.2 show more varied patterns as seen on Figures 4(a) and 4(b). When injecting the Native American task vector (Figure 12), accuracy remains stable while fairness seems to decrease (increased DPD and EOD). Asian (Figure 11) shows the same behavior as injecting the Men task vector (Figure 5(a)), positive increase of fairness metrics as $\lambda$ increases. These results show that injecting task vectors shifts fairness and performance in a group-specific manner, tracing a clear fairness–utility frontier. This heterogeneity is expected: per §5.2 and Theorem Theorem (informal), sensitivity scales with $|\Delta\theta_{g}|_{2}$. Practically, task vector merging thus offers a subgroup-conditioned control knob: identifying which $\Delta\theta_{g}$ help or hurt which groups provides a new actionable design consideration that SFT/LoRA do not expose, and that hasn’t been explored in previous task arithmetic literature.
@@ Conclusion.
In this study, we investigated how task arithmetic with task vectors affects group fairness, in comparison to conventional FFT and LoRA methods. We conducted experiments to assess how adding a fairness-oriented task vector impacts prediction accuracy and fairness metrics, including DPD and EOD across multiple subgroups. The results indicate that, with appropriate settings of the scalar coefficient $\lambda$, task arithmetic can improve DPD and EOD without significantly compromising overall model accuracy. Notably, low to moderate values of $\lambda$ often reduced prediction bias in minority groups compared to FFT and LoRA, a pattern we observe across three datasets (hate speech, toxicity, and age detection) and four model families (LLaMA‑2, DistilBERT, Qwen‑2.5, ViT-Base).
Furthermore, the task-arithmetic formulation exposes a single interpretable parameter $\lambda$, which allows practitioners to trace subgroup-specific changes along a fairness–accuracy trade-off curve. This transparency makes it easier to diagnose when edits disproportionately benefit or harm particular groups.
@@
Despite these promising results, several limitations remain. First, our experiments focus on the open-weight 0.5–7B regime (DistilBERT, Qwen-0.5B, LLaMA-2-7B, ViT-Base), where task arithmetic is practically deployable and comparable to prior work. This setting gives us full weight access to construct subgroup task vectors and compute exact DPD/EOD. We do not evaluate against proprietary frontier models; prior work shows that prompt-based debiasing can offer partial gains but is brittle and does not provide principled control over group-level disparities (Ma et al., 2023; Furniturewala et al., 2024; Chu et al., 2024). Extending our framework to larger or API-only models is an important direction for future work. Second, we study a simple scaling scheme based on a shared global coefficient $\lambda$. This keeps the intervention interpretable but cannot capture richer multidimensional fairness constraints. More expressive parameterizations (e.g., multiple coefficients or optimization over combinations of $\lambda$, represent natural extensions). Finally, our evaluation covers a limited set of classification tasks with single sensitive attributes. The effectiveness of task arithmetic may depend on dataset and subgroup structure, so assessing generalization to broader domains and intersectional settings remains important.
Reproducibility statement.
We provide code  [^5] Our code is available at https://github.com/LauraGomezjurado/fairness_task_vector_deploy, configs, and scripts to reproduce all experiments, including data preprocessing, training, and evaluation. All datasets and base models used are open-source/publicly available; we include scripts to fetch the exact versions. Exact hyperparameters, model identifiers, and implementation details are documented in the appendix, along with seeds and hardware/software specs. Results are reported over multiple runs, and we provide instructions to regenerate all figures and tables from logged outputs.
Acknowledgments
Our deepest gratitude goes out to the anonymous reviewers whose invaluable insights substantially enhanced the quality of this manuscript. This work was supported by RBC Borealis through the RBC Borealis AI Global Fellowship Award, which was awarded to Hiroki Naganuma. We are also profoundly grateful to The Masason Foundation for their generous support in providing computational resources and for fostering an environment that encourages deep research collaboration. The computation resources were further supported by “TSUBAME Encouragement Program for Young/Female Users” of Center for Information Infrastructure at Institute of Science Tokyo and by “Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures” in Japan.
Appendix
Appendix A Extended Related Works
Task arithmetic: efficiency and interpretability.
Task vectors offer a computationally efficient framework for editing and analyzing model behavior. Once a task vector is computed—namely, the weight difference between a base model and its fine-tuned variant (Ilharco et al., 2023; Zhang et al., 2024; Yoshida et al., 2025b)—no additional training data or retraining is required to transfer or remove task-specific capabilities. By treating each fine-tuning update as a direction in weight space, practitioners can combine or negate these updates through simple addition or subtraction (Ilharco et al., 2023). This modularity not only reduces computational overhead but also enhances interpretability by isolating the contribution of each task. Beyond modularity, task arithmetic can reveal valuable information about how and where a model adapts to new tasks. Li et al. (2024) show a near-linear relationship between data size and the norm of a task vector, suggesting that over-represented tasks can dominate weight space shifts in multi-task settings. In addition, the orientation of task vectors can indicate synergies or conflicts among tasks (Li et al., 2025), and decomposing these vectors by layer can pinpoint which parts of the model are most affected (Zhang et al., 2024; Gargiulo et al., 2025). Hence, task vectors offer a promising lens for diagnosing training dynamics and identifying potential biases.
Vision–Language Models and Demographic Bias Audits.
Audits of vision–language systems consistently report demographic biases in zero-shot classification, retrieval, and captioning, along with mixed effectiveness of post-hoc debiasing methods (ali2023evaluated). Large studies of CLIP-style models show that scaling data or parameters does not reliably reduce these disparities, which vary with dataset composition and training regimes (Sahili et al., 2025). Even current vision–language models exhibit measurable group-level gaps under standard prompting (Wu et al., 2025).
Appendix B Fairness metrics
B.1 Demographic Parity Difference (DPD) (Agarwal et al., 2018; 2019)